Eric spunkel is three times as old as she is now 4 years ago he was four times ashold as she was then how old is Eric now
Answer by ankor@dixie-net.com(22740)

let e = Eric's age now
let h = her age now
:
Eric spunkel is three times as old as she is now.
e = 3h
4 years ago he was four times as old as she was then
e - 4 = 4(h-4)
e - 4 = 4h - 16
e = 4h - 16 + 4
e = 4h - 12
replace e with 3h
3h = 4h - 12
12 = 4h - 3h
h = 12 yr is her age now
;
how old is Eric now
3(12) = 36 yrs